OpenStudy (anonymous):

FInd the exact value of csc (21 degrees)

OpenStudy (anonymous):

csc(21)

OpenStudy (anonymous):

how do i input this into the calculator?

OpenStudy (owlcoffee):

A very important trigonometric operation, being this, the inverse of sine is defined as the "cosecant" of any angle, writing in a mathematcal standard: \[\frac{ 1 }{ \sin v }=\csc v\]

zepdrix (zepdrix):

Is the problem really asking for "exact value"? Because calculator probably won't help with that.

OpenStudy (anonymous):

You have \(\csc \theta=\dfrac{1}{\sin \theta}\), so let's work with \(\sin\) for the moment. Fair warning: I prefer using radians over degrees - too many \({}^\circ\) `\circ` symbols to keep track of with the latter. Hopefully you know that \(21^\circ=\dfrac{\pi}{60}\text{ rad}\). Notice that\[\frac{\pi}{60}=\frac{\pi}{10}-\frac{\pi}{12}\]which means \[\sin\frac{\pi}{60}=\sin\left(\frac{\pi}{10}-\frac{\pi}{12}\right)=\sin\frac{\pi}{10}\cos\frac{\pi}{12}-\sin\frac{\pi}{12}\cos\frac{\pi}{10}\]Easy enough, but not sufficient if you don't know the trig ratio values of \(\dfrac{\pi}{10}\) and \(\dfrac{\pi}{12}\). Let's derive them. \(\dfrac{\pi}{12}\) is much simpler to work with. Recalling the half-angle identity, you have that \[\sin^2\frac{\pi}{12}=\frac{1}{2}\left(1-\cos\frac{\pi}{6}\right)\implies\sin\frac{\pi}{12}=\frac{\displaystyle \sqrt{2-\sqrt{3}}}{2}\\ \cos^2\frac{\pi}{12}=\frac{1}{2}\left(1+\cos\frac{\pi}{6}\right)\implies\cos\frac{\pi}{12}=\frac{\displaystyle \sqrt{2+\sqrt{3}}}{2}\]Not too challenging, I hope? The next part is a bit more involved. Again making use of the half-angle identity, you get \[\begin{cases}x=\sin\dfrac{\pi}{10}\\[1ex] y=\cos\dfrac{\pi}{10}\end{cases} \implies \cos\frac{\pi}{5}=\begin{cases}1-2x^2\\[1ex] 2y^2-1\end{cases}\]Recall Euler's formula, \[e^{it}=\cos t+i\sin t\]and DeMoivre's theorem, which states that for non-negative integers \(n\), \[(\cos t+i\sin t)^n=\left(e^{it}\right)^n=e^{int}=\cos nt+i\sin nt\]Let \(n=5\) (why will be apparent in just a moment), then \[\cos5t+i\sin 5t=(\cos t+i\sin t)^5\]Expanding and equating the real parts, you have \[\cos5t=\cos^5t-10\cos^3t\sin^2t+5\cos t\sin^4t\]Via the Pythagorean identity, we can almost make this equation a polynomial in \(\cos\): \[\cos5t=\cos^5t-10\cos^3t\left(1-\cos^2t\right)+5\left(1-\cos^2t\right)^2\cos t\]Notice that if we set \(t=\dfrac{\pi}{5}\), then we end up with an equation quintic in \(\cos\dfrac{\pi}{5}\). Doing so, expanding, and setting \(z=\cos\dfrac{\pi}{5}\), yields \[-1=16z^5-20z^3+5z\]This equation is actually quite easy to solve. Observe that it has a root at \(z=-1\), so \(z+1\) is a factor. \[\begin{align*} 0&=16z^5-20z^3+5z+1\\[1ex] &=(z+1)(4z^2-2z-1)^2\\[1ex] z&=-1,\,\frac{1\pm\sqrt5}{4} \end{align*}\]Since \(\dfrac{\pi}{5}\) is an angle in the first quadrant (ie. \(0<\dfrac{\pi}{5}<\dfrac{\pi}{2}\)), it's obvious that \(z=\cos\dfrac{\pi}{5}\) is positive, which means \(z=\dfrac{1+\sqrt5}{4}\). So, you find that \[\cos\frac{\pi}{5}=\frac{1+\sqrt5}{4}=\begin{cases}1-2x^2\\[1ex]2y^2-1\end{cases} \implies \begin{cases} x=\sin\dfrac{\pi}{10}=\displaystyle\sqrt{\frac{1-\dfrac{1+\sqrt5}{4}}{2}}\\[1ex] y=\cos\dfrac{\pi}{10}=\displaystyle\sqrt{\frac{\dfrac{1+\sqrt5}{4}+1}{2}} \end{cases}\] Finally (!), you arrive at an exact form: \[\begin{align*}\sin\frac{\pi}{60}&=\left(\displaystyle\sqrt{\frac{1-\dfrac{1+\sqrt5}{4}}{2}}\right)\left(\frac{\displaystyle \sqrt{2+\sqrt{3}}}{2}\right)-\left(\frac{\displaystyle \sqrt{2-\sqrt{3}}}{2}\right)\left(\displaystyle\sqrt{\frac{\dfrac{1+\sqrt5}{4}+1}{2}}\right)\\[2ex] &=\frac{\sqrt{2(2+\sqrt3)(3-\sqrt5)}-\sqrt{2(2-\sqrt3)(5+\sqrt5)}}{8} \end{align*}\]which in turn yields \[\csc21^\circ=\csc\frac{\pi}{60}=\frac{8}{\sqrt{2(2+\sqrt3)(3-\sqrt5)}-\sqrt{2(2-\sqrt3)(5+\sqrt5)}}\]